OpenStudy (anonymous):

The area of a circle A is given by the formula A = ∏ r^2, where r is the radius of the circle. Find the radius of a circle with area 16 in.^2. Round to the nearest tenth of an inch.

OpenStudy (campbell_st):

well you have \[16 = \pi r^2\] so its \[r^2 = \frac{16}{\pi}\] just take the square root for the answer

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Ok thank you.